OpenStudy (anonymous):

I'm working with exponents and I have a fraction of 6 to the 40 power over 6 to the 10 power. How do I do it? Help!!!!

OpenStudy (anonymous):

Is it???? 6^40 ----- 6^10

OpenStudy (anonymous):

If yes, then using the Exponential Law 6^40 ----- = 6^40-10 6^10

OpenStudy (anonymous):

If yes, then using the Exponential Law 6^40 ----- = 6^40-10 = 6^30 6^10

jhonyy9 (jhonyy9):

so this is 6^(40-10) =6^30

OpenStudy (anonymous):

We use the law \[a ^{m}/a ^{n} = a ^{m-n}\]